### v1

# Mechanochemical feedback drives complex inertial dynamics in active solids

Abstract
Active solids combine internal active driving with elasticity to realize states with nonequilibrium mechanics and autonomous motion. They are often studied in overdamped settings, e.g., in soft materials, and the role of inertia is less explored. We construct a model of a chemically active solid that incorporates mechanochemical feedback and show that, when feedback overwhelms mechanical damping, autonomous inertial dynamics can spontaneously emerge through sustained consumption of chemical fuel. By combining numerical simulations, analysis and dynamical systems approaches, we show how active feedback drives complex nonlinear dynamics on multiple time-scales, including limit cycles and chaos. Our results suggest design principles for creating ultrafast actuators and autonomous machines from soft, chemically-powered solids.
Introduction. Soft materials driven by chemical stimuli have been widely used as actuators to achieve complex tasks—locomotion, grasping, and shape-morphing—through flexible deformations [1, 2]. While promising for technologies [3], their contraction speed and power output are often limited by slow, diffusive mechanisms (thermal, fluidic, ionic, etc.) [4, 5] and damping. To overcome this limitation, recent approaches have used miniaturization [6, 7] or inertial dynamics triggered by plant-inspired elastic instabilities [8, 9, 10] and fast chemical processes (e.g., combustion) [11, 12, 13, 14] to achieve fast actuation. However, such systems passively relax without sustained input, prompting us to ask: How can soft, chemically driven materials sustain fast and continuous inertial motion without cyclic stimulation, and how complex can this motion be?
A potential strategy is suggested by contractile tissues that use biological feedback to drive extreme motions in animals [15, 16]. Such active solids, exemplified by living tissues [17, 18, 19, 16], cellular and organismal collectives [20, 21], mechanochemical gels [22, 23, 24] and robotic assemblies [25], have been shown to produce exotic dynamics [26, 27, 19, 28, 29, 18, 20, 16] and present new opportunities for designing soft chemically-driven machines.
Building upon previous studies on biological systems with overdamped feedback [27, 30, 31, 19, 32, 16, 33] and inertial robotic assemblies [34, 35, 36], we combine both effects to demonstrate how chemically active soft solids with sufficiently strong mechanochemical feedback can sustain complex and autonomous nonlinear inertial dynamics. We show how feedback introduces new timescales that govern how stress accelerates or decelerates chemical reactions. When this response becomes faster than viscous dissipation of momentum, the system continually extracts chemical energy to power self-sustained inertial oscillations and chaotic dynamics. This suggests mechanochemical feedback as a promising scalable mechanism for devising soft engines, actuators, and autonomous machines with fast and rich dynamics.
Mechanochemical active solid. We model the active solid as a one-dimensional (1D) gel of length $L$ and mass density $\rho$ suffused with a mixture of chemical species $A$ and $B$ (concentration $n_{A},n_{B}$) that inter-convert into each other via a reversible reaction $A\xrightleftharpoons{}B$. The total concentration is conserved by the reaction ($n_{A}+n_{B}=n_{0}$), so the relative concentration difference defined by $\chi=(n_{B}-n_{A})/n_{0}$ serves as a convenient reaction coordinate. Material deformations are assumed small and captured by a linearized strain $\epsilon(x,t)=\partial_{x}u(x,t)$ ($u$: displacement, $x$: material coordinate). The free energy of the system [37, 38, 39, 40],
$\mathcal{F}=k_{B}T\int\mathrm{d}x\left[n_{0}f(\chi)+\dfrac{E}{2k_{B}T}\epsilon^{2}+n_{0}C\chi\epsilon\right]\;,$ (1)
combines a local free energy density $f(\chi)$, an elastic energy with modulus $E$ and a passive coupling $C$ between the chemical concentration and deformations, with $T$ the temperature. The passive coupling controls the chemically induced mechanical stresses present, e.g., as can be seen in chemically responsive gels [22, 24], solvent-driven swelling [9], and explosives or combustion [14, 41].
The mechanical stress in the material $\sigma=\sigma^{p}+\sigma^{d}$ consists of both passive ($\sigma^{p}=\delta\mathcal{F}/\delta\epsilon=E\epsilon+n_{0}k_{B}TC\chi$) and dissipative ($\sigma^{d}=\eta\partial_{t}\epsilon$) contributions [38, 39, 40], with $\eta$ a viscosity. In the absence of any reactive chemicals ($n_{0}=0$), the material behaves as a passive Kelvin-Voigt viscoelastic solid. Kinematics and momentum conservation then gives
$\partial_{t}u=v\;,$
$\rho\partial_{t}v=\partial_{x}(\sigma^{p}+\sigma^{d})\;,$ (2)
where $v(x,t)$ is the local material velocity.
Chemical reactions involve transitions between molecular states and, in equilibrium, a balance of forward and backward transition rates [42, 43]. That mechanical deformations can modify reaction rates (kinetics) and bias chemical equilibria (energetics) is known in diverse systems, e.g., mechanochemically responsive polymers [44, 45, 46], inorganic compunds [47, 48], biomolecules [49, 50, 51], and living cells [52, 53].
Following standard transition state theory [42, 54] generalized for non-ideal systems [55, 56], we have forward and backward reaction rates given by $r_{+}=k_{0}\exp(\mu_{A}/k_{B}T)$ and $r_{-}=k_{0}\exp(\mu_{B}/k_{B}T)$, respectively, where $\mu_{A},\mu_{B}$ are the chemical potentials. Here $k_{0}$ is a microscopic rate constant that depends on details of the transition state. In ideal systems, these rates reduce to the law of mass action. In a passive system, $\mu^{p}_{i}=\delta\mathcal{F}/\delta n_{i}$, so $\mu^{p}_{B}=-\mu_{A}^{p}=k_{B}T[f^{\prime}(\chi)+C\epsilon]$. This relation demonstrates reciprocity between the reaction and strain ($\delta\mu_{i}^{p}/\delta\epsilon=\delta\sigma^{p}/\delta n_{i}=\mp k_{B}TC$, $i=A,B$) [37, 43] and captures the energetic shift in chemical equilibria due to deformations.
To drive the system out of chemical equilibrium, we augment the reaction rates by setting $\mu_{i}=\mu_{i}^{p}+\mu^{a}_{i}$, where an external source of reactive free energy, say from an abundant supply of a fuel that catalyses the reaction [39, 40], drives a bias in chemical potential. A constant bias simply changes the equilibrium concentration of reactants and products by a fixed amount and leaves the system passive. But a dynamic bias can induce an active coupling between mechanical stress and the reaction beyond the free energy in Eq. (1). In particular, mechanical stress can modify both the transition state barrier and the chemical potential difference, which we minimally capture by setting $\mu^{a}_{i}=h_{i}\sigma$, where $\sigma$ is the total stress and $h_{i}$ are mechanochemical coupling constants. This feedback breaks reciprocity as $\delta\mu_{i}/\delta\epsilon\neq\delta\sigma^{p}/\delta n_{i}$, and provides a nonequilibrium coupling between the reaction and mechanical degrees of freedom.
Collecting these effects together, and assuming $\mu_{i}\ll k_{B}T$ (the chemical potential associated with the reaction is small) to linearize the reaction rates, we obtain (see more details in the Supplementary Materials (SM) [57])
$\partial_{t}\chi=r_{+}-r_{-}=r_{\mathrm{eq}}+\zeta\sigma,$ (3)
where $r_{\mathrm{e}q}=(k_{0}/k_{B}T)(\mu_{A}-\mu_{B})=-2k_{0}[f^{\prime}(\chi)+C\epsilon]$ is the equilibrium reaction rate including the passive coupling to strain and $\zeta=k_{0}(h_{A}-h_{B})/k_{B}T$ is the non-equilibrium coupling to stress. Eqs. (Mechanochemical feedback drives complex inertial dynamics in active solids -3) along with initial and boundary conditions completely specify the inertial active solid model.
Figure 1: Mechanochemical coupling. (a,b) Feedback loops among reaction ($\chi$), strain ($\partial_{x}u$), strain rate ($\partial_{x}v$) in equilibrium (a) and active (b) solids. (c) Contour plot of reaction rate $\partial_{t}\chi$ (in A.U.) using the 0D model in Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids ), with the dashed line marking $\partial_{t}\chi=0$ so $\chi$ grows (decays) above (below) this curve. An example trajectory is shown where a perturbation from the origin leads to a limit cycle, with four representative states illustrated in the insets.
Figure 2: Dynamics of the 1D mechanochemical system in Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids ). We set $L=2\pi$, $\eta=0.16$, and $\tau_{c}=100$ along with (a) $\alpha_{u}=50,\alpha_{v}=0$, (b) $\alpha_{u}=-10,\alpha_{v}=0$, (c) $\alpha_{u}=-50,\alpha_{v}=0$, (d) $\alpha_{u}=50,\alpha_{v}=77$, (e) $\alpha_{u}=50,\alpha_{v}=80$, (f) $\alpha_{u}=50,\alpha_{v}=82$. We include a small diffusion constant $D=0.05$ to regularize chemical gradients. The initial conditions are $u(x,0)=0.01\sin(x)$, $v(x,0)=0=\chi(x,0)$. The first row shows the displacement field $u(x,t)$. The second row shows the phase space $(\partial_{x}u,v)$ trajectory at the point $x=\pi/2$.
1D model. To exemplify its key consequences, we consider a bistable chemical free energy $f(\chi)=-a_{2}\chi^{2}/2+a_{4}\chi^{4}/4$ with $a_{2},a_{4}>0$. The model (Eqs. (Mechanochemical feedback drives complex inertial dynamics in active solids -3)) then simplifies to give
$\partial_{t}u=v\;,$
$\rho\partial_{t}v=E\partial_{x}^{2}u+\eta\partial_{x}^{2}v+c\partial_{x}\chi,$ (4)
$\tau_{c}\partial_{t}\chi=\chi(1-g\chi^{2})-\alpha_{u}\partial_{x}u-\alpha_{v}\partial_{x}v,$
where $c=n_{0}k_{B}TC$ is the passive mechanochemical coupling, $\tau_{c}=1/[2k_{0}a_{2}+\zeta c]$ is the normalized chemical reaction timescale, $g=2k_{0}a_{4}\tau_{c}$ controls the chemical nonlinearity, and $\alpha_{u}=\tau_{c}(2k_{0}C-\zeta E)$ and $\alpha_{v}=-\tau_{c}\zeta\eta$ are the nonequilibrium feedback couplings to strain and strain-rate, respectively (Fig. 1a,b).
There are three homogeneous steady-state solutions, the trivial one ($\chi=\chi_{ss},\epsilon=\epsilon_{ss}$, and $v=0$), which always exists, and two non-trivial ones ($\epsilon_{ss}=\pm(c/E)\sqrt{(E+c\alpha_{u})/Eg}$, $\chi_{ss}=\mp\sqrt{(E+c\alpha_{u})/Eg}$, and $v=0$), which only exist for $c\alpha_{u}<-E$. We numerically solve Eqs. (Mechanochemical feedback drives complex inertial dynamics in active solids ) in a 1D domain of length $L=2\pi$ with periodic boundary conditions using the pseudo-spectral method [58]. We pick time, length, mass, and reaction coordinate scales so that $\rho=E=c=g=1$ and vary the activity parameters, $\alpha_{u}$ and $\alpha_{v}$.
In Fig. 2, we showcase rich dynamics that can emerge from this 1D model. For $\alpha_{v}=0$ (only strain feeding back to reaction), as $\alpha_{u}$ decreases from positive (reciprocal) to negative (nonreciprocal), the system changes from a phase separated steady state (Fig. 2a) to a homogeneous steady state when $\alpha_{u}<-E/c$ (Fig. 2b), and develops a limit cycle (LC) at large negative $\alpha_{u}$ (Fig. 2c). On the other hand, fixing $\alpha_{u}>0$ and increasing $\alpha_{v}$ (strain rate also feeds back to reaction) triggers nonequilibrium states with increasing complexity: LC (Fig. 2d), period doubling (Fig. 2e), and chaos (Fig. 2f). These dynamical states caused by $\alpha_{u}<0$ or $\alpha_{v}>0$ are sustained by continuous conversion of chemical energy to mechanical energy. While nonreciprocal couplings are known to be a generic mechanism for creating dynamic patterns in systems with conservation laws [26, 59, 60, 61], here we focus on their interplay with inertia.
Scaling analysis reveals the key features underlying the transitions between the qualitative behaviors exhibited in Fig. 2. In the absence of a chemical reaction ($\chi=0$), long-wavelength deformations on the scale of the system size $L$ transport momentum by elastic waves on an inertial timescale $\tau_{\mathrm{el}}\sim L\sqrt{\rho/E}$, and dissipate it through viscous stresses on a diffusive (Stokes) timescale $\tau_{\eta}\sim\rho L^{2}/\eta$. Importantly, we note that $\tau_{\eta}$ depends on inertia and is distinct from the local viscoelastic relaxation time $\tau_{ve}\sim\eta/E$. Reaction kinetics on the other hand proceeds on a chemical timescale $\sim\tau_{c}$, but nonequilibrium mechanochemical feedback can pump energy into the system at different rates. To see this, we note that a chemically induced stress $\sim c\chi$ generates a characteristic elastic strain $|\epsilon|\sim c\chi/E$ and viscous strain rate $|\dot{\epsilon}|\sim c\chi/\eta$ that drive the reaction on two new time-scales, $\tau_{1}\sim\tau_{c}E/|\alpha_{u}c|$ and $\tau_{2}\sim\tau_{c}\eta/|\alpha_{v}c|$, set by the strain and strain-rate feedback terms respectively (Fig. 1b). When activity overwhelms the rate of mechanical dissipation, i.e., $\tau_{1},\tau_{2}\lesssim\tau_{\eta},\tau_{ve}$, we expect complex dynamics to emerge (Fig. 2). Note that for the examples in Fig. 2 we choose $\eta=0.16$, so $\tau_{\eta}>\tau_{\mathrm{el}}$ and the system is mechanically underdamped.
Figure 3: Dynamical phases due to nonreciprocal strain-chemistry feedback ($A_{v}=0,A_{u}<0$) in the 0D model (Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids )). (a) Phase diagram ($\mathsf{Ch}=100$). (b-d) Typical trajectories in the $(U,V)$ plane for (b) the LC with $A_{u}=-11.0$, (c) one FP at $U_{0}=X_{0}=0,V_{0}=0$ with $A_{u}=-6.0$, and (d) FPs with the trajectory converging to one of them $U_{+}=-X_{+}=\sqrt{1+A_{u}}$ with $A_{u}=-0.5$. The FP is shown as red dots in (b,c), and red (unstable) and green (stable) dots in (d). We use $\mathsf{Vi}=0.10$ in all cases. (e) Dependence of the time period $T$ on $A_{u}$ for different $\mathsf{Ch}$ in the large LC region (in $\tau_{\mathrm{el}}=1$ units).
Figure 4: Dynamical phases due to nonreciprocal strain-strain rate-chemistry feedback ($A_{v}>0,A_{u}>0$) in the 0D model (Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids )). (a) Phase diagram ($\mathrm{Ch}=100$, $A_{u}=50$). (b-f) Typical trajectories in the $(U,V)$ plane for the FPs with the trajectory converging to (b) one FP $U_{+}=-X_{+}=\sqrt{1+A_{u}}$ for $A_{v}=60.5$, (c) an LC around one unstable FP for $A_{v}=68.0$, (d) period doubling for $A_{v}=78.5$, (e) chaos for $A_{v}=84.0$, and (f) the large LC conclusing all FPs for $A_{v}=106.0$. In all cases, we set $\mathsf{Vi}=0.16$ and $A_{u}=50.0$. (g) Bifurcation diagram showing the route from LC around the FP $U_{\pm}=-X_{\pm}=\sqrt{1+A_{u}}$ to chaos as a function of the parameter $A_{v}$. Here, $u_{\mathrm{m}}$ represents the local maximum value of $U$. (h) Dependence of the time period $T$ on $A_{v}$ at $\mathsf{Vi}=0.16$ in the large LC region, approaching $2\pi$ as $A_{v}$ increases (in $\tau_{\mathrm{el}}=1$ units).
Effective 0D model. We next perform a single mode Galerkin approximation to reduce the spatial complexity of the 1D model by projecting Eq. (Mechanochemical feedback drives complex inertial dynamics in active solids ) onto the first fundamental mode of the 1D system. We set $u(x,t)=(1/q)U(t)\sin(qx)$, $v(x,t)=(\sqrt{E/\rho})V(t)\sin(qx)$, and $\chi(x,t)=(E/c)X(t)\cos(qx)$, with $q=2\pi/L$ to non-dimensionalize the equations and obtain
$\tau_{\mathrm{el}}\;\dot{U}=V\;,$
$\tau_{\mathrm{el}}\;\dot{V}=-U-\mathsf{Vi}\;V-X\;,$ (5)
$\tau_{c}\;\dot{X}=X(1-X^{2})-A_{u}\;U-A_{v}\;V\;,$
where we have assumed $g=(4/3)(c/E)^{2}$ to simplify the chemical nonlinearity and set $\tau_{\mathrm{el}}=(L/2\pi)\sqrt{\rho/E}$ as the inertial time for a long-wavelength elastic wave. The relative importance of viscous dissipation to inertia defines a visco-inertial number $\mathsf{Vi}=2\pi\eta/(L\sqrt{\rho E})\sim\tau_{\mathrm{el}}/\tau_{\eta}$. Similarly, the ratio of the chemical reaction time to that of inertia defines a chemo-inertial number $\mathsf{Ch}=(2\pi\tau_{c}/L)\sqrt{E/\rho}=\tau_{c}/\tau_{\mathrm{el}}$. As we are primarily interested in the mechanically underdamped regime with slow chemical reactions, we expect $\tau_{\mathrm{el}}<\tau_{\eta}<\tau_{c}$, i.e., $\mathsf{Vi}<1$ and $\mathsf{Ch}\gg 1$. Finally, nonequilibrium activity is captured by two feedback parameters: $A_{u}=\alpha_{u}c/E\sim\tau_{c}/\tau_{1}$ and $A_{v}=2\pi\alpha_{v}c/(L\sqrt{\rho E})\sim\mathsf{Vi}\;\tau_{c}/\tau_{2}$. Note that when $A_{u}<0$, the mechanochemical coupling is nonreciprocal. With initial conditions, the four nondimensional parameters $\mathsf{Vi}$, $\mathsf{Ch}$, $A_{u}$, and $A_{v}$ completely define the response of our 0D system. Mathematically similar models appear in quantum and classical control systems [62, 63, 64, 65, 66, 67], metamaterials [68] and in the physics of hearing [69, 70], but the characterization of the nonlinear dynamics in the context of inertial mechanochemical solids is novel to our knowledge. As in the 1D model, for steady-state, Eqs. (Mechanochemical feedback drives complex inertial dynamics in active solids ) has (i) one fixed point (FP) $U_{0}=X_{0}=0$ for $A_{u}<-1$, and (ii) three FPs $U_{\pm}=-X_{\pm}=\pm\sqrt{1+A_{u}}$ and $U_{0}=X_{0}=0$ for $A_{u}>-1$ (at FPs, by definition, $V=0$).
At $A_{v}=0$, nonreciprocity between strain and reaction arises for $A_{u}<0$. Upon decreasing $A_{u}$, a linear stability analysis reveals first a pitchfork bifurcation where 3 FPs (bistable) transition to 1 FP (monostable) and then an oscillatory Hopf bifurcation occurs at a second threshold when $A_{u}\approx-\mathsf{Vi}\;\mathsf{Ch}$ [71] (Fig. 3). The analytical bifurcation thresholds (derived in the SM [57]) are confirmed via numerical continuation using MATCONT and shown in Fig. 3a. This Hopf bifurcation corresponds to $\tau_{1}\lesssim\tau_{\eta}$, i.e., when the strain feedback actively pumps energy faster than viscous stresses can dissipate it, and it crucially depends on inertia as $\tau_{\eta}\sim\rho L^{2}/\eta$.
Near the Hopf bifurcation, the imaginary part of the unstable eigenvalue yields the time period of the LC to be $T=2\pi\tau_{\mathrm{el}}/\sqrt{1-(\mathsf{Vi}/\mathsf{Ch})}\simeq 2\pi\tau_{\mathrm{el}}$ for $\mathsf{Ch}\gg 1$ (dashed line in Fig. 3e): fast oscillations at the inertial time scale, distinct from previous examples of spontaneous oscillations in overdamped active solids [72, 19, 27, 16, 17]. Interestingly, this instability and LC behavior persists even in the mechanically overdamped regime ($\mathsf{Vi}>1$, Fig. 3) with a period controlled by the same $\tau_{el}$, as long as nonreciprocal feedback overwhelms the mechanical dissipation rate (see more details in the SM [57]).
Now we consider $A_{u}>0,A_{v}>0$, where strain rate steers the activity (Fig. 4). By choosing $A_{u}>0$, we always have three FPs, with the FP at the origin ($U_{0},X_{0}$) always unstable. A linear stability analysis shows that upon increasing $A_{v}$, both nontrivial FPs $(U_{\pm},X_{\pm})$ lose stability via a Hopf bifurcation when $A_{v}\gtrsim\mathsf{Ch}$ (valid in the limit of slow kinetics, i.e., $\mathsf{Ch}\gg\mathsf{Vi},A_{u}$, see SM [57] for the full expression). This analytical bifurcation threshold (white dashed line in Fig. 4(a)) is confirmed by numerical analysis. In terms of physical time scales, this bifurcation condition corresponds to $\tau_{2}<\tau_{ve}$, i.e., the strain rate feedback time is smaller than the local viscoelastic relaxation time. Near onset, the small LC has a slow oscillation time period $T_{\mathrm{small}}\approx\pi\tau_{\mathrm{el}}\sqrt{2\mathsf{Vi}\,\mathsf{Ch}/(1+A_{u})}\sim\sqrt{\eta\tau_{c}/E}$ governed by dissipative processes. Note that these timescales do not rely on either spatial gradients or inertia, and the resulting instability is smoothly connected to previously studied instabilities in overdamped active solids [72, 19, 27, 16, 17].
At larger $A_{v}$, more complex dynamics arises, enabled by the three-dimensional phase space ($U,V,X$) including inertia. We find a period doubling cascade for the pair of LCs, which collide into a chaotic strange attractor (Fig. 4e), reminiscent of the classical Lorenz and Rossler system [73]. Chaos develops in a large parameter regime in the $A_{v}-\mathsf{Vi}$ space (Fig. 4a), and is interspersed by small islands of periodic orbits (Fig. 4g and SM [57] for Lyapunov exponent analysis). The phase boundaries separating chaos from limit cycles in Fig. 4(a) are obtained by analyzing phase-space trajectories obtained by direct numerical simulations. For even larger values of $A_{v}$, eventually, the system settles into a globally attracting large LC that encloses all three FPs (Fig. 4f). While the small LCs have slow oscillations, we numerically find that the large LC exhibits fast inertial oscillations with a time period approaching $T\approx 2\pi\tau_{\mathrm{el}}$ as $A_{v}\to\infty$ (Fig. 4h). This explains and recapitulates the complex behavior observed in the 1D model (Fig. 2).
Conclusion. By using nonequilibrium feedback to couple chemical reactions with the mechanics of a soft solid, we have shown how active driving can sustain autonomous deformations, fast oscillations, and nonlinear dynamics. These results highlight the importance of inertia in active mechanochemical solids and provide a design principle for generating high quality ultrafast oscillators even in dissipative, soft solids. That feedback can also trigger chaos in active solids suggests a simple way to generate complex “twitching” movements, reminiscent of rudimentary behaviors in soft-bodied organisms [74, 15]. Our work paves the way to explore a new regime of chemically active solids in previously inaccessible high frequency scales, and provides a scalable and efficient platform to design autonomous soft materials from homogeneous media, without the need for complex microstructures or discretely embedded active particles.
Acknowledgments.–N.B., X.M. and B.A. acknowledge support from the US Army Research Office (Grant No. W911NF-20-2-0182). X.M., S.S. and Y.W. acknowledge support from the Office of Naval Research (MURI N00014-20-1-2479). X.M. acknowledge support from National Science Foundation through the Materials Research Science and Engineering Center at the University of Michigan, Award No. DMR-2309029. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).

### v2

@@
 Soft materials driven by chemical stimuli have been widely used as actuators to achieve complex tasks—locomotion, grasping, and shape-morphing—through flexible deformations [1, 2].
-While promising for technologies [3], their contraction speed and power output are often limited by slow, diffusive mechanisms (thermal, fluidic, ionic, etc.) [4, 5] and damping.
+While promising for technologies [3], their contraction speed and power output are often limited by slow, diffusive mechanisms (thermal, fluidic, ionic, etc.) and damping [4, 5].
 To overcome this limitation, recent approaches have used miniaturization [6, 7] or inertial dynamics triggered by plant-inspired elastic instabilities [8, 9, 10] and fast chemical processes (e.g., combustion) [11, 12, 13, 14] to achieve fast actuation.
@@
 $\partial_{t}\chi=r_{+}-r_{-}=r_{\mathrm{eq}}+\zeta\sigma,$ (3)
-where $r_{\mathrm{e}q}=(k_{0}/k_{B}T)(\mu_{A}-\mu_{B})=-2k_{0}[f^{\prime}(\chi)+C\epsilon]$ is the equilibrium reaction rate including the passive coupling to strain and $\zeta=k_{0}(h_{A}-h_{B})/k_{B}T$ is the non-equilibrium coupling to stress.
+where $r_{\mathrm{e}q}=(k_{0}/k_{B}T){(\mu^{p}_{A}-\mu^{p}_{B})}=-2k_{0}[f^{\prime}(\chi)+C\epsilon]$ is the equilibrium reaction rate including the passive coupling to strain and $\zeta=k_{0}(h_{A}-h_{B})/k_{B}T$ is the non-equilibrium coupling to stress.
+Note that for the system to be passive, we need $\zeta=0$, which invalidates coupling to strain rate.
 Eqs.
@@
 Figure 1: Mechanochemical coupling.
-(a,b) Feedback loops among reaction ($\chi$), strain ($\partial_{x}u$), strain rate ($\partial_{x}v$) in equilibrium (a) and active (b) solids.
+(a,b) Feedback loops among reaction ($\chi$), strain ($\partial_{x}u$), strain rate ($\partial_{x}v$) in passive (a) and active (b) solids.
+In the passive limit (a), detailed balance enforces vanishing strain-rate coupling, and a reciprocal strain-reaction coupling ($k_{B}TC$).
+When active (b), the feedback drives nonequilibrium fluxes on independent time scales $\tau_{1,2}$.
 (c) Contour plot of reaction rate $\partial_{t}\chi$ (in A.U.) using the 0D model in Eq.
 (Mechanochemical feedback drives complex inertial dynamics in active solids), with the dashed line marking $\partial_{t}\chi=0$ so $\chi$ grows (decays) above (below) this curve.
-An example trajectory is shown where a perturbation from the origin leads to a limit cycle, with four representative states illustrated in the insets.
+An example trajectory is shown where a perturbation from the origin leads to a limit cycle, with four representative states illustrated as insets, where the background shading refers to concentration $\chi$.
+Starting with the bottom left corner, increased strain breaks stress-sensitive crosslinkers, which drives $\chi$ to increase, followed by decrease in strain, and recovery back to the first state.
 Figure 2: Dynamics of the 1D mechanochemical system in Eq.
@@
 We include a small diffusion constant $D=0.05$ to regularize chemical gradients.
-The initial conditions are $u(x,0)=0.01\sin(x)$, $v(x,0)=0=\chi(x,0)$.
-The first row shows the displacement field $u(x,t)$.
-The second row shows the phase space $(\partial_{x}u,v)$ trajectory at the point $x=\pi/2$.
+The initial conditions are $u(x,0)=0.01\sin(x)$, $v(x,0)=0=\chi(x,0)$ The first row shows the displacement field $u(x,t)$.
+The second row shows the phase space $({\epsilon=\partial_{x}u},v)$ trajectory at the point $x=9\pi/25$.
 1D model.
@@
 $\tau_{c}\partial_{t}\chi=\chi(1-g\chi^{2})-\alpha_{u}\partial_{x}u-\alpha_{v}\partial_{x}v,$
-where $c=n_{0}k_{B}TC$ is the passive mechanochemical coupling, $\tau_{c}=1/[2k_{0}a_{2}+\zeta c]$ is the normalized chemical reaction timescale, $g=2k_{0}a_{4}\tau_{c}$ controls the chemical nonlinearity, and $\alpha_{u}=\tau_{c}(2k_{0}C-\zeta E)$ and $\alpha_{v}=-\tau_{c}\zeta\eta$ are the nonequilibrium feedback couplings to strain and strain-rate, respectively (Fig.
+where $c=n_{0}k_{B}TC$ is the passive chemically induced stress scale, $\tau_{c}=1/[2k_{0}a_{2}+\zeta c]$ is the effective chemical reaction timescale, combining the equilibrium rate ($2k_{0}a_{2}$) and the active feedback rate ($\zeta c$[57], $g=2k_{0}a_{4}\tau_{c}$ controls the chemical nonlinearity, and $\alpha_{u}=\tau_{c}(2k_{0}C-\zeta E)$ and $\alpha_{v}=-\tau_{c}\zeta\eta$ are the nonequilibrium feedback couplings to strain and strain-rate, respectively (Fig.
 1a,b).
@@
 Importantly, we note that $\tau_{\eta}$ depends on inertia and is distinct from the local viscoelastic relaxation time $\tau_{ve}\sim\eta/E$.
-Reaction kinetics on the other hand proceeds on a chemical timescale $\sim\tau_{c}$, but nonequilibrium mechanochemical feedback can pump energy into the system at different rates.
-To see this, we note that a chemically induced stress $\sim c\chi$ generates a characteristic elastic strain $|\epsilon|\sim c\chi/E$ and viscous strain rate $|\dot{\epsilon}|\sim c\chi/\eta$ that drive the reaction on two new time-scales, $\tau_{1}\sim\tau_{c}E/|\alpha_{u}c|$ and $\tau_{2}\sim\tau_{c}\eta/|\alpha_{v}c|$, set by the strain and strain-rate feedback terms respectively (Fig.
+The bare reaction kinetics on the other hand proceeds on a chemical timescale $\sim\tau_{c}$, but nonequilibrium mechanochemical feedback can pump energy into the system at independent rates.
+To see this, we consider the time taken for a minimal feedback cycle (Fig.
+1b,c).
+A chemically induced stress $\sim c\chi$ generates a characteristic elastic strain $|\epsilon|\sim c\chi/E$ or viscous strain rate $|\dot{\epsilon}|\sim c\chi/\eta$ that can in turn drive the reaction via $\tau_{c}\dot{\chi}\sim\alpha_{u}\epsilon$ or $\tau_{c}\dot{\chi}\sim\alpha_{v}|\dot{\epsilon}|$, completing the cycle.
+This gives two new time-scales, $\tau_{1}\sim\tau_{c}E/|\alpha_{u}c|$ and $\tau_{2}\sim\tau_{c}\eta/|\alpha_{v}c|$, set by the strain and strain-rate feedback terms respectively (Fig.
 1b).
@@
 (a) Phase diagram ($\mathsf{Ch}=100$).
-(b-d) Typical trajectories in the $(U,V)$ plane for (b) the LC with $A_{u}=-11.0$, (c) one FP at $U_{0}=X_{0}=0,V_{0}=0$ with $A_{u}=-6.0$, and (d) FPs with the trajectory converging to one of them $U_{+}=-X_{+}=\sqrt{1+A_{u}}$ with $A_{u}=-0.5$.
-The FP is shown as red dots in (b,c), and red (unstable) and green (stable) dots in (d).
-We use $\mathsf{Vi}=0.10$ in all cases.
-(e) Dependence of the time period $T$ on $A_{u}$ for different $\mathsf{Ch}$ in the large LC region (in $\tau_{\mathrm{el}}=1$ units).
+(b-d) Typical trajectories in the $(U,V)$ plane for (b) the outer (larger) LC with $\left(\mathsf{Vi},A_{u}\right)=\left(0.70,-130.0\right)$ (underdamped regime) and the inner (smaller) LC with $\left(\mathsf{Vi},A_{u}\right)=\left(1.2,-140.0\right)$ (overdamped regime), (c) one FP at $U_{0}=X_{0}=0,V_{0}=0$ with $\left(\mathsf{Vi},A_{u}\right)=\left(0.70,-20.0\right)$, and (d) FPs with the trajectory converging to one of them $U_{+}=-X_{+}=\sqrt{1+A_{u}}$ with $\left(\mathsf{Vi},A_{u}\right)=\left(0.70,-0.5\right)$.
+The FPs are shown as red (unstable) and green (stable) dots in (b-d), with the arbitrarily chosen initial conditions denoted by the black star.
+In panels (b-d), different symbols adjacent to the panel index represent the corresponding points in the phase diagram in (a).
+(e) Dependence of the time period $T$ on $A_{u}$ for different $\mathsf{Ch}$ in the LC region (in $\tau_{\mathrm{el}}=1$ units).
 Figure 4: Dynamical phases due to nonreciprocal strain-strain rate-chemistry feedback ($A_{v}>0,A_{u}>0$) in the 0D model (Eq.
@@
 (a) Phase diagram ($\mathrm{Ch}=100$, $A_{u}=50$).
-(b-f) Typical trajectories in the $(U,V)$ plane for the FPs with the trajectory converging to (b) one FP $U_{+}=-X_{+}=\sqrt{1+A_{u}}$ for $A_{v}=60.5$, (c) an LC around one unstable FP for $A_{v}=68.0$, (d) period doubling for $A_{v}=78.5$, (e) chaos for $A_{v}=84.0$, and (f) the large LC conclusing all FPs for $A_{v}=106.0$.
+(b-f) Typical trajectories in the $(U,V)$ plane for the FPs with the trajectory converging to (b) one FP $U_{+}=-X_{+}=\sqrt{1+A_{u}}$ for $A_{v}=60.5$, (c) an LC around one unstable FP for $A_{v}=68.0$, (d) period doubling for $A_{v}=78.5$, (e) chaos for $A_{v}=84.0$, and (f) the large LC enclosing all FPs for $A_{v}=106.0$.
 In all cases, we set $\mathsf{Vi}=0.16$ and $A_{u}=50.0$.
+The black stars in (b-f) denote the arbitrarily chosen initial condition.
 (g) Bifurcation diagram showing the route from LC around the FP $U_{\pm}=-X_{\pm}=\sqrt{1+A_{u}}$ to chaos as a function of the parameter $A_{v}$.
-Here, $u_{\mathrm{m}}$ represents the local maximum value of $U$.
+Here, $u_{\mathrm{m}}$ represents the local maximum value of $U$ [57].
 (h) Dependence of the time period $T$ on $A_{v}$ at $\mathsf{Vi}=0.16$ in the large LC region, approaching $2\pi$ as $A_{v}$ increases (in $\tau_{\mathrm{el}}=1$ units).
@@
 $\tau_{\mathrm{el}}\;\dot{V}=-U-\mathsf{Vi}\;V-X\;,$ (5)
-$\tau_{c}\;\dot{X}=X(1-X^{2})-A_{u}\;U-A_{v}\;V\;,$
+${\tau_{\mathrm{el}}\mathsf{Ch}}\;\dot{X}=X(1-X^{2})-A_{u}\;U-A_{v}\;V\;,$
 where we have assumed $g=(4/3)(c/E)^{2}$ to simplify the chemical nonlinearity and set $\tau_{\mathrm{el}}=(L/2\pi)\sqrt{\rho/E}$ as the inertial time for a long-wavelength elastic wave.
@@
 3).
+Figure 3 shows typical trajectories in the $(U,V)$ plane for limit cycles (b) and FPs (c-d).
 The analytical bifurcation thresholds (derived in the SM [57]) are confirmed via numerical continuation using MATCONT and shown in Fig.
@@
 3e): fast oscillations at the inertial time scale, distinct from previous examples of spontaneous oscillations in overdamped active solids [72, 19, 27, 16, 17].
-Interestingly, this instability and LC behavior persists even in the mechanically overdamped regime ($\mathsf{Vi}>1$, Fig.
-3) with a period controlled by the same $\tau_{el}$, as long as nonreciprocal feedback overwhelms the mechanical dissipation rate (see more details in the SM [57]).
+Interestingly, this instability and LC behavior persists even in the mechanically overdamped regime ($\mathsf{Vi}>1$, see SM [57] for more details) with a period controlled by the same $\tau_{\mathrm{el}}$, as long as nonreciprocal feedback overwhelms the mechanical dissipation rate (Fig.
+3b,e).
 Now we consider $A_{u}>0,A_{v}>0$, where strain rate steers the activity (Fig.
 4).
-By choosing $A_{u}>0$, we always have three FPs, with the FP at the origin ($U_{0},X_{0}$) always unstable.
+By choosing $A_{u}>0$, we have three FPs, with the FP at the origin ($U_{0},X_{0}$) always unstable.
 A linear stability analysis shows that upon increasing $A_{v}$, both nontrivial FPs $(U_{\pm},X_{\pm})$ lose stability via a Hopf bifurcation when $A_{v}\gtrsim\mathsf{Ch}$ (valid in the limit of slow kinetics, i.e., $\mathsf{Ch}\gg\mathsf{Vi},A_{u}$, see SM [57] for the full expression).
@@
 Near onset, the small LC has a slow oscillation time period $T_{\mathrm{small}}\approx\pi\tau_{\mathrm{el}}\sqrt{2\mathsf{Vi}\,\mathsf{Ch}/(1+A_{u})}\sim\sqrt{\eta\tau_{c}/E}$ governed by dissipative processes.
-Note that these timescales do not rely on either spatial gradients or inertia, and the resulting instability is smoothly connected to previously studied instabilities in overdamped active solids [72, 19, 27, 16, 17].
-At larger $A_{v}$, more complex dynamics arises, enabled by the three-dimensional phase space ($U,V,X$) including inertia.
-We find a period doubling cascade for the pair of LCs, which collide into a chaotic strange attractor (Fig.
+Note that these timescales do not rely on inertia, and the resulting instability is smoothly connected to previously studied instabilities in overdamped active solids [72, 19, 27, 16, 17].
+Enabled by the three-dimensionality of phase space ($U,V,X$) including inertia, at large $A_{v}$, chaos arises.
+We find a period doubling (Fig.
+4d) cascade for the pair of LCs, which collide into a chaotic strange attractor (Fig.
 4e), reminiscent of the classical Lorenz and Rossler system [73].
@@
 2).
+Whether such chaotic and fast dynamics observed for $A_{v}>0$ survives when $\mathsf{Vi}\gtrsim 1$ is left for future work.
 Conclusion.
-By using nonequilibrium feedback to couple chemical reactions with the mechanics of a soft solid, we have shown how active driving can sustain autonomous deformations, fast oscillations, and nonlinear dynamics.
+By using nonequilibrium feedback to couple chemical reactions with the mechanics of a soft solid, we have shown how active driving can sustain autonomous deformations, fast oscillations, and complex dynamics.
 These results highlight the importance of inertia in active mechanochemical solids and provide a design principle for generating high quality ultrafast oscillators even in dissipative, soft solids.
@@
 Our work paves the way to explore a new regime of chemically active solids in previously inaccessible high frequency scales, and provides a scalable and efficient platform to design autonomous soft materials from homogeneous media, without the need for complex microstructures or discretely embedded active particles.
+[…]
 Acknowledgments.–N.B., X.M.